OpenStudy (winsteria):

FAN AND MEDAL Seth compares the two new scenarios and concludes that driving from a relative’s house is faster than biking from Dad’s work. Bike from Work: His dad’s work is 5 miles from school, and he bikes at 0.2 mile a minute. Relative’s House: y = 1.2x − 4 Seth learns that there is construction on the highway from his relative’s house, slowing the rate to 0.7 mile each minute. How much longer will it take from his relative’s house during the construction compared to biking?

OpenStudy (winsteria):

@baru @AaronAndyson

OpenStudy (winsteria):

@robtobey

OpenStudy (winsteria):

@jango_IN_DTOWN

OpenStudy (winsteria):

anyone know?

OpenStudy (bibby):

the question seems incomplete to me. I mean if dads work is 5 miles away and he bikes at 0.2miles a minute 5 mile = 0.2mi*xminutes \(\dfrac{5}{0.2}=xminutes\\x=25 minutes\) so we know it takes 25 minutes to get there by bike, I don't know what to do with the 2nd half

OpenStudy (winsteria):

oh well that's the question :/

OpenStudy (winsteria):

It wants to know how much minuets slower

OpenStudy (bibby):

uhh maybe they want something like y=0.2x y=0.7x -4 how much slower is the first than the last \((0.7x-4)-0.2x=0.5x-4\) now plug in x=25 \(12.5-4=8.5\) I'm not positive on this

OpenStudy (winsteria):

well It's okay thank you anyway maybe you could help me with the last one

OpenStudy (bibby):

sure, post it